# Multi-Class Stackelberg Games for the Co-Design of Networked Systems

### v1

@@ Abstract
We investigate a co-design problem, encompassing simultaneous design of system infrastructure and control, through a game-theoretical framework. To this end, we propose the co-design problem as a two-layer hierarchical strategic interaction. At the upper layer, a leader (or multiple leaders) determines system design parameters, while at the lower layer, a follower (or multiple followers) optimizes the control strategy. To capture this hierarchy, we propose four novel classes of Stackelberg games that integrate diverse strategic behaviors, including combinations of cooperative and non-cooperative interactions across two different layers. Notably, the leaders’ interactions are represented using a normal-form game, whereas the followers’ interactions are modeled by different games (dynamic games in discrete time). These distinct game structures result in a Stackelberg game that accommodates different game types per layer, and/or supports heterogeneous strategic behaviors involving cooperation and non-cooperation simultaneously. Learning algorithms using the best-response dynamics are used to solve the game problems when considering a discrete strategic space for the leaders. The efficacy of the proposed approach is demonstrated through an application to the co-design of the Barcelona drinking water network.
@@ I Introduction
Game theory allows the study and modeling of strategic interactions, and it has been extensively used in the solution of engineering problems, e.g., for power systems, robot coordination, water distribution systems, etc. Several classes of games have been studied and reported in the literature, capturing different situations, such as cooperation, competition, altruism, and co-opetition. One of the most well-known game-theoretical solution concepts is the Nash equilibrium, which describes a situation in which there are no incentives for a single decision-maker to modify his/her selection in an unilateral way. This solution concept takes place in the non-cooperative games [1]. However, more possibilities have been explored in the game theory community, e.g., Stackelberg games to describe major and minor decision-makers or hierarchical interactions [2], [3, Chapter 7], coalitional games to understand the power index and/or level of influence within a strategic interaction [4, 5], cooperative games describing common interest payoffs, Berge games to describe altruism [6], co-opetition to describe simultaneously levels of cooperation and competition [7], [3, Chapter 5], among others. Indeed, it has been extensively shown that game theory is a suitable and powerful theoretical tool to model non-centralized control strategies.
@@
Modern networked systems are inherently complex and large-scale, comprising multiple interconnected subsystems that must coordinate to achieve overall system efficiency and reliability [8, 9]. Examples of such systems include water distribution networks [10, 11], power grids [12], supply chains [13], and transportation systems [14]. The design and control of such systems pose significant challenges, particularly in industrial applications where system infrastructure is typically designed based on conservative estimates, and advanced control strategies are developed afterward. This sequential approach often results in suboptimal performance, as the control strategy is constrained by a fixed system design. To address this issue, we focus on co-design problem, an approach that simultaneously optimizes both system infrastructure and the associated control strategy [15]. A solution of the co-design problem could achieve the global optimum regarding to the overall objectives.
In this work, we focus on a co-design problem, which is suitably modeled by considering the combination of game-theoretical solution concepts at different stages. From a global view, we consider two layers, which can be interpreted as the designer of the system and the entity in charge of its operation. Yet, we can study strategic interaction possibilities within each one of these layers. We pursue the study of how designer decisions can be modeled when multiple entities are making decisions at this level, or we might consider that this layer is composed of a unique entity performing the design and solving an optimization problem. Regarding the operator of the system, we might encounter multiple options for strategic behavior that mainly span centralized and non-centralized approaches.
@@
The main contribution of this paper is to propose a two-layer hierarchical game or a Stackelberg-like game problem that combines at each layer different classes of games. For the leaders’ layer, we have a normal form non-cooperative game with a finite number of strategies and whose utility functions depend on both the leaders’ and followers’ strategic selection. Hence, at this leaders’ layer, we may also consider cooperation, in which we perform optimization for the design parameters of the system. Regarding the followers’ layer, we introduce a dynamic game, more precisely, a non-cooperative difference game problem. Similarly, we may also consider a cooperative game at the followers’ layer, which corresponds to a standard optimal control problem [16]. We highlight the fact that the interaction between the layers in the hierarchical scheme is different from others reported in the literature. The hierarchical scheme coupling is mainly given by the fact that the leader defines the feasible set of strategies for the follower. We show that the proposed Stackelberg game, which combines leader normal form games with discrete strategic sets and follower difference games (dynamic games), can be used for solving the co-design control and system problem. The four Stackelberg game classes that we study in this paper are:
@@ II Problem Statement
This paper considers a networked engineering system and aims to analyze two main problems using game-theoretical tools. On one hand, we analyze the design of the networked system led by a leader’s layer. At this design stage, some key parameters for the system are determined. On the other hand, we analyze the control design to operate the system led by a followers’ layer. At this design stage, optimal control policies are defined to operate the system subject to the established designed parameters by the leaders. There are two classes of decision-makers, i.e., leaders and followers. The set of $L\in\mathbb{N}$ leaders is given by :=$\mathcal{L}:=\{1,\dots,L\}$, and the set of $M\in\mathbb{N}$ followers is given by :=$\mathcal{M}:=\{1,\dots,M\}$. The leader decision-makers set takes care of the system design and parameter setting, whereas the second decision-makers set is in charge of designing the optimal control to operate the system, adhering to the imposed rules of the leader’s set. The leaders interact with each other using a static normal form game, and their decisions are coupled with the followers’ interactions. The followers interact with each other by means of a difference game whose settings are leader-strategic-dependent.
@@ II-B Followers Non-Cooperative Behavior
Once the follower’s game problem is defined, we introduce the equilibrium concept by using the best-response strategies. First, we present the best way a follower can react to a design in Definition 1.
@@
Then, we use the best-response strategy concept to introduce the game-theoretical equilibrium concept in Definition 2 below.
@@
A strategic profile $\mathbf{u}^{*}=(\mathbf{u}^{*}_{1},\dots,\mathbf{u}^{*}_{M})\in\mathcal{U}$ is a Nash equilibrium if all the strategies are best-response strategies against each other given the strategic selection of the leader(s), i.e., $\mathbf{u}^{*}_{i}\in\mathbb{BR}^{\mathcal{M}}_{i}(\mathbf{u}^{*}_{-i},\{\text{leader strategies}\})$, for all $i\in\mathcal{M}$. The set of Nash equilibria for the followers strategic interaction is denoted by $\mathbb{NE}_{\mathrm{follower}}$. $\square$
Later on, in Section , we develop some algorithms to find the Nash equilibria. Under these algorithms, one might arbitrarily converge to a point close to the Nash equilibrium. Thus, Definition 3 below introduces the concept of $\varepsilon$ equilibrium, in which we can finish the solution seeking.
@@
We also would like to study the case in which the followers interact in a cooperative framework. This strategic behavior is formally presented in the next subsection.
@@ II-C Followers Cooperative Behavior
Similar to the Assumption 1 stated in the non-cooperative game problem settings, we state the following assumption for the cooperative case.
@@
Next, we introduce the best action that a follower can take, in a cooperative way, against the decisions made by the leader, in Definition 5. Note that this is different from the Definition 1.
@@
We denote $\mathbb{BR}^{\mathrm{C}}_{\mathcal{M}}(\{\text{leader strategies}\})$ as the set of cooperative best response strategies given the strategic selection by the leader. $\square$
Once the strategic interactions for the followers are introduced, we proceed to formally present the strategic interactions corresponding to the leaders.
@@ II-D Leaders Strategic Interaction and Settings
The leaders action set is $\{A_{1},\dots,A_{L}\}$ and each set contains $N\in\mathbb{N}$ actions. The $j-$th leader chooses an action $a_{j}\in A_{j}$, and a strategic profile is given by the joint selection of strategies along all the players, i.e.,
@@ II-E Leaders Non-Cooperative Behavior
There are two important aspects to mention related to the game settings for the leaders’ problem. On one hand, note that the strategic set is considered to the discrete, i.e., there is a finite number of possible decisions that the leaders can make. On the other hand, we are only evaluating the pure strategies for the leaders. Therefore, we need to set the following assumption.
@@
To complete all the possible combinations in the interactions among leaders and followers, we also consider the case in which the leaders behave in a cooperative way as introduced in the following subsection.
@@ II-G Price of Anarchy for Leaders and Followers
respectively. A lower price of anarchy corresponds to an enhancement in the worst scenario for the equilibrium performance, and $\mathrm{PoA}_{\mathrm{leader}}=1$ or $\mathrm{PoA}_{\mathrm{follower}}=1$ imply that the Nash equilibria are optimal. In other words, when having a unitary price of anarchy, the non-centralized design by the leaders using normal form games, or the non-centralized design of the control law by the followers’ means of dynamic games, are as optimal as those obtained when they jointly optimize in the framework of cooperation.
III Stackelberg equilibrium for
the co-design problem
@@ III Stackelberg equilibrium for
Figure 1: Different hierarchical strategic interactions illustrating the four Stackelberg game classes (Squares represent leaders and circles represent followers).
As we can consider multiple strategic interactions at each layer of the hierarchical Stackelberg scheme, we have multiple classes for the Stackelberg game as summarized in Table I. Each class is also presented in Fig. 1.
@@ III-A Definition of Stackelberg Equilibrium Classes
Let us now assume that all the leaders $\mathcal{L}$ coordinate each other and agree to jointly perform the design of the system. This cooperation leads to an optimization that is jointly performed by the $L$ leaders. Then, once the system design is established, the followers $\mathcal{M}$ cooperatively design the control actions to operate the system. This cooperative followers’ behavior leads to a centralized control design. Definition 11 formally presents the Stackelberg equilibrium under this scenario (see Fig. 1).
@@
In this case, all the leaders jointly design parameters against which the followers will optimally respond also jointly by means of cooperation. $\square$
The aforementioned Stackelberg game classes have considered a homogeneous behavior for both the leader and followers, i.e. both layers cooperating or both layers acting independently in a non-cooperative framework. However, layers may exhibit heterogeneous strategic behavior. First, let us consider the case in which the leaders $\mathcal{L}$ do not cooperate for the system design, whereas the followers $\mathcal{M}$ cooperatively react to the system design by solving jointly a centralized optimal control problem. The emerging Stackelberg equilibrium for this scenario is formally presented next in Definition 12 (see Fig. 1).
@@
The Stackelberg equilibria for the different classes imply the computation of either a Nash equilibrium given by a fixed point condition or the solution of an optimization problem for the followers in terms of the leaders, and then the same procedure for the leader knowing the optimal response of the followers against the leaders’ decisions. In the following Section, algorithms to solve each one of the classes are presented.
@@ III-B Learning Procedure to Compute Stackelberg Solutions
This section formally presents the procedure to compute each one of the Stackelberg equilibria corresponding to the four different classes introduced above (see Fig. 1). We present a qualitative description of the steps that comprise each one of the computations together with its corresponding algorithm.
@@ III-B1 Class I: Stackelberg game with non-cooperative leaders and non-cooperative followers
Given that under this strategic behavior combination, both sets of decision-makers, i.e., leaders $\mathcal{L}$ and $\mathcal{M}$, play in a non-cooperative manner, the solution for the normal form game and dynamic game problems can be computed by following best-response dynamics. Then, the Stackelberg equilibrium is found for this game class. The steps are the following:
  • For each feasible leader strategic selection $\mathbf{a}\in\mathcal{A}$, let us compute the Nash equilibrium $\mathbf{u}^{*}\in\mathcal{U}$, which is denoted by $\mathbf{u}^{*}(\mathbf{a})$, meaning that the equilibrium strategic profile depends on the strategic selection of the leader. The computed Nash equilibrium is a non-cooperative best response against the leaders’ strategies, i.e., $\mathbf{u}^{*}(\mathbf{a})\in\mathbb{BR}^{\mathrm{NC}}_{\mathcal{M}}(\mathbf{a})$.
  • Using the equilibrium strategic profile $\mathbf{u}^{*}(\mathbf{a})$ in terms of the leaders’ selections, the leader non-cooperative game problem can be stated only in terms of $\mathbf{a}$, and the Nash equilibrium for the leaders $\mathbf{a}^{*}\in\mathcal{A}$ can be found.
  • Then, using the Nash equilibrium for the leaders game problem $\mathbf{a}^{*}$, the Nash equilibrium for the follower game problem becomes $\mathbf{u}^{*}:=\mathbf{u}^{*}(\mathbf{a}^{*})$.
  • Finally, one obtains the Class I Stackelberg equilibrium as $(\mathbf{u}^{*},\mathbf{a}^{*})\in\mathcal{U}\times\mathcal{A}$.

### v2

@@ Abstract
We investigate a co-design problem, which entails the simultaneous optimization of both system infrastructure and control strategies, within a game-theoretical framework. To this end, we formulate the co-design problem as a two-layer hierarchical strategic interaction. At the upper layer, a leader (or multiple leaders) determines system design parameters, while at the lower layer, a follower (or multiple followers) optimizes the control strategy. To capture this hierarchy, we propose four novel classes of Stackelberg games that integrate diverse strategic behaviors, including combinations of cooperative and non-cooperative interactions across two different layers. Notably, the leaders’ interactions are represented using a normal-form game, whereas the followers’ interactions are modeled by difference games (dynamic games in discrete time). These distinct game structures result in a Stackelberg game formulation that accommodates different game types per layer, and/or supports heterogeneous strategic behaviors involving cooperation and non-cooperation simultaneously. Learning algorithms using the best-response dynamics are used to solve the game problems when considering a discrete strategic space for the leaders. The effectiveness of the proposed approach is demonstrated through an application to the co-design of the Barcelona drinking water network.
@@ I Introduction
Game theory provides a rigorous mathematical framework for analyzing and modeling strategic interactions among rational decision-makers. Its versatility has led to widespread applications in engineering, including power systems [1], robotic coordination [2], and cyber-physical systems [3], among others. The literature identifies several classes of games that capture distinct interaction dynamics, such as cooperation, competition, altruism, and co-opetition.
Several classes of games have been extensively studied and reported in the literature, each tailored to represent specific interactive scenarios. One of the most fundamental and widely recognized solution concepts in non-cooperative game theory is the Nash equilibrium, which characterizes a stable outcome wherein no agent has an incentive to unilaterally deviate from its current strategy [4]. Beyond this, the game theory community has developed a variety of other frameworks to capture more nuanced interaction structures. Stackelberg game is a hierarchical decision-making process, where one player (the leader) moves first, and the others (the followers) react accordingly [5], [6, Chapter 7]. Coalitional games have been developed to analyze the formation of alliances, compute power indices, and assess the influence of individual participants in cooperative environments [7]. Similarly, cooperative games explore settings in which agents seek to maximize a collective payoff, while Berge games are used to formalize altruistic behaviors in strategic contexts [8]. The concept of co-opetition, which blends elements of cooperation and competition, has been formalized to describe many real-world scenarios where agents must balance conflicting objectives [9], [6, Chapter 5]. Given this diversity of models, it is well-established that game theory provides a powerful theoretical tool for designing and analyzing decentralized control strategies in complex systems.
@@
Modern networked systems are inherently complex and large-scale, comprising multiple interconnected subsystems that must coordinate to achieve overall system efficiency and reliability [10, 11]. Examples of such systems include water distribution networks [12, 13], power grids [14], supply chains [15], and transportation systems [16]. The design and control of these systems pose significant challenges, particularly in industrial applications where infrastructure is often developed based on conservative estimates, with advanced control strategies introduced only afterward. This sequential design–then–control approach leads to a suboptimal solution, as the control strategy is constrained by the limitations of a predetermined system design. To address this limitation, we focus on co-design problem, an approach that seeks to simultaneously optimize both the system infrastructure and the associated control strategy [17]. By jointly considering these two aspects, the co-design framework has the potential to achieve globally optimal solutions with respect to the overall system objectives, thereby overcoming the inherent inefficiencies of sequential design.
In this paper, we study the co-design problem based on Stackelberg game. The Stackelberg game framework, originally proposed by Heinrich von Stackelberg in [5], models hierarchical decision-making structures involving a leader and one or more followers. The use of leader-follower schemes has been successfully applied to solve several engineering problems, e.g., formation control [18, 19], crowd dynamics modeling [20], hierarchical scheme for price dynamics analysis under the framework of mean-field-type games [21]. In its original formulation, Stackelberg analyzed the interactions between firms with asymmetric decision-making, where the leader commits to a strategy first, and the followers respond optimally to that strategy. This hierarchical, bi-level decision-making framework is well-suited for addressing the co-design problem, as it mirrors the natural structure of infrastructure design (leader) followed by operational control (follower). Accordingly, we adopt a Stackelberg game formulation to model the inter-dependencies between system design and control, thereby providing a structured methodology for deriving optimal co-design solutions.
Furthermore, we extend this modeling framework by considering the possibility of strategic interactions within each layer of the hierarchical system. From a global perspective, we conceptualize the problem as consisting of two interacting layers: the designer of the system and the operator responsible for its ongoing functionality. Within the design layer, we may encounter multiple decision-makers with potentially conflicting objectives, or alternatively, a centralized designer solving an optimization problem on behalf of the entire system. Similarly, on the operational side, we may face either centralized or decentralized control strategies, depending on the system characteristics and operational constraints. This layered and strategic view provides a flexible and comprehensive framework for modeling complex co-design problems in large-scale networked systems.
@@
The main contribution of this paper is to propose a two-layer hierarchical game or a Stackelberg-like game problem that combines at each layer different classes of games. For the leaders’ layer, we have a normal form non-cooperative game with a finite number of strategies and whose utility functions depend on both the leaders’ and followers’ strategic selection. Hence, at this leaders’ layer, we may also consider cooperation, in which we perform optimization for the design parameters of the system. Regarding the followers’ layer, we introduce a dynamic game, more precisely, a non-cooperative difference game problem. Similarly, we may also consider a cooperative game at the followers’ layer, which corresponds to a standard optimal control problem [22]. We highlight that the interaction between the layers in the hierarchical scheme is different from others reported in the literature. The hierarchical scheme of coupling is mainly given by the fact that the leader defines the feasible set of strategies for followers.
We show that the proposed Stackelberg game, which combines leader normal form games with discrete strategic sets and follower difference games (dynamic games), can be used for solving the co-design control and system problem. The four Stackelberg game classes that we study in this paper are:
@@ II Problem Statement
In this paper, we consider a networked engineering system and aim to analyze two main problems using game-theoretical tools. On one hand, we analyze the design of the networked system led by a leader’s layer. At this design stage, some key parameters for the system are determined. On the other hand, we analyze the control design to operate the system led by a followers’ layer. At this design stage, optimal control policies are defined to operate the system subject to the established design parameters by the leaders. There are two classes of decision-makers, i.e., leaders and followers. The set of $L\in\mathbb{N}$ leaders is given by :=$\mathcal{L}:=\{1,\dots,L\}$, and the set of $M\in\mathbb{N}$ followers is given by :=$\mathcal{M}:=\{1,\dots,M\}$. The leader decision-makers set takes care of the system design and parameter setting, whereas the second decision-makers set is in charge of designing the optimal control to operate the system, adhering to the imposed rules of the leader’s set. The leaders interact with each other using a static normal form game, and their decisions are coupled with the followers’ interactions. The followers interact with each other by means of a difference game whose settings are leader-strategic-dependent.
@@ II-B Followers Non-Cooperative Behavior
Once the follower’s game problem is defined, we introduce the equilibrium concept by using the best-response strategies.
@@
Then, we use the best-response strategy concept to introduce the definition of game-theoretical equilibrium below.
@@
A strategic profile $\mathbf{u}^{*}=(\mathbf{u}^{*}_{1},\dots,\mathbf{u}^{*}_{M})\in\mathcal{U}$ is a Nash equilibrium if all the strategies are best-response strategies against each other given the strategic selection of the leader(s), i.e., $\mathbf{u}^{*}_{i}\in\mathbb{BR}^{\mathcal{M}}_{i}(\mathbf{u}^{*}_{-i},\{\text{leader strategies}\})$, for all $i\in\mathcal{M}$. The set of Nash equilibria for the followers’ strategic interaction is denoted by $\mathbb{NE}_{\mathrm{follower}}$. $\square$
The following definition introduces the concept of $\varepsilon$ equilibrium, in which we can finish the solution seeking.
@@
We would also like to study the case in which the followers interact in a cooperative framework. This strategic behavior is formally presented in the next subsection.
@@ II-C Followers Cooperative Behavior
Similar to Assumption 1 stated in the non-cooperative game problem settings, the following assumption is made for the cooperative case.
@@
Next, we introduce the best action that a follower can take, in a cooperative way, against the decisions made by the leader, in Definition 5. Note that this is different from Definition 1.
@@
We denote $\mathbb{BR}^{\mathrm{C}}_{\mathcal{M}}(\{\text{leader strategies}\})$ as the set of cooperative best response strategies given the strategic selection by the leaders. $\square$
@@ II-D Leaders Strategic Interaction and Settings
The leaders action set is $\{A_{1},\dots,A_{L}\}$ and each set contains $N\in\mathbb{N}$ actions. The $j-$th leader chooses an action $a_{j}\in A_{j}$. A strategic profile is given by the joint selection of strategies along all the players, i.e.,
@@ II-E Leaders Non-Cooperative Behavior
There are two important aspects to mention related to the game settings for the leaders’ problem. On one hand, note that the strategic set is considered to be discrete, i.e., a finite number of possible decisions that the leaders can make. On the other hand, we are only evaluating the pure strategies for the leaders. Therefore, we need to set the following assumption.
@@
To complete all the possible combinations in the interactions among leaders and followers, we also consider the case in which the leaders behave in a cooperative way, as introduced in the following subsection.
@@ II-G Price of Anarchy for Leaders and Followers
respectively. A lower price of anarchy corresponds to an enhancement in the worst scenario for the equilibrium performance, and $\mathrm{PoA}_{\mathrm{leader}}=1$ or $\mathrm{PoA}_{\mathrm{follower}}=1$ imply that the Nash equilibria are optimal. In other words, when having a unitary price of anarchy, the non-centralized design by the leaders using normal form games, or the non-centralized design of the control law by the dynamic games of followers, are as optimal as those obtained when they jointly optimize in the framework of cooperation.
@@
III Stackelberg equilibrium for
the co-design problem
As we can consider multiple strategic interactions at each layer of the hierarchical Stackelberg scheme, we have multiple classes for the Stackelberg game as summarized in Table I. Each class is also presented in Fig. 1. It is worth highlighting that, unlike other works employing Stackelberg games, where strategic coupling between the leader and the follower typically occurs through either the dynamics of the system or cost/utility functional (see [6, 24, 25]), the formulation considered in this work focuses on a different type of interaction. Specifically, we analyze the case in which the leader’s decision affects the feasible set of the optimization problem of the followers. This distinction allows us to capture a broader class of hierarchical decision-making scenarios, particularly relevant for co-design problems where infrastructure choices constrain the subsequent control strategies.
Figure 1: The hierarchical strategic interactions illustrating four Stackelberg game classes (Squares represent multiple leaders and circles represent multiple followers).
@@ III-A Definition of Stackelberg Equilibrium Classes
Let us now assume that all the leaders $\mathcal{L}$ coordinate with each other and agree to jointly perform the design of the system. This cooperation leads to an optimization that is jointly performed by the $L$ leaders. Then, once the system design is established, the followers $\mathcal{M}$ cooperatively design the control actions to operate the system. This cooperative followers’ behavior leads to a centralized control design. Definition 11 formally presents the Stackelberg equilibrium under this scenario (see Fig. 1).
@@
In this case, all the leaders jointly design parameters against which the followers will optimally respond, also jointly by means of cooperation. $\square$
The aforementioned Stackelberg game classes have considered a homogeneous behavior for both the leader and followers, i.e., both layers cooperating or both layers acting independently in a non-cooperative framework. However, layers may exhibit heterogeneous strategic behavior. First, let us consider the case in which the leaders $\mathcal{L}$ do not cooperate for the system design, whereas the followers $\mathcal{M}$ cooperatively react to the system design by solving jointly a centralized optimal control problem. The emerging Stackelberg equilibrium for this scenario is formally presented next in Definition 12.
@@
The Stackelberg equilibria for the different classes imply the computation of either a Nash equilibrium given by a fixed point condition or the solution of an optimization problem for the followers in terms of the leaders, and then the same procedure for the leader, knowing the optimal response of the followers against the leaders’ decisions. In the following Section, algorithms to solve each of the classes are presented.
@@ III-B Learning Procedure to Compute Stackelberg Solutions
This section provides a formal description of the procedures required to compute the Stackelberg equilibria for the four classes of games introduced in Fig. 1. For each case, we present a qualitative overview of the computational steps along with the associated algorithmic implementation. We use the following notation: $|A_{j}|$ is the cardinality of the strategic set $A_{j}$ and $A_{j}(s)$ denotes the $s$-th element from $A_{j}$.
@@ III-B1 Class I: Stackelberg game with non-cooperative leaders and non-cooperative followers
Given that under this strategic behavior combination, both sets of decision-makers, i.e., leaders $\mathcal{L}$ and $\mathcal{M}$, play in a non-cooperative manner, the solution for the normal form game and dynamic game problems can be computed by following best-response dynamics. In this configuration, the leaders select their strategies independently and selfishly, anticipating the equilibrium response of the followers, who also act independently in optimizing their individual objectives.
The procedure for computing the Stackelberg equilibrium in this context involves the following steps:
  • Follower Response: For each feasible leader strategic selection $\mathbf{a}\in\mathcal{A}$, let us compute the Nash equilibrium $\mathbf{u}^{*}\in\mathcal{U}$, which is denoted by $\mathbf{u}^{*}(\mathbf{a})$, meaning that the equilibrium strategic profile depends on the strategic selection of the leader. The computed Nash equilibrium is a non-cooperative best response against the leaders’ strategies, i.e., $\mathbf{u}^{*}(\mathbf{a})\in\mathbb{BR}^{\mathrm{NC}}_{\mathcal{M}}(\mathbf{a})$.
  • Leader Game Solution: Using the equilibrium strategic profile $\mathbf{u}^{*}(\mathbf{a})$ in terms of the leaders’ selections, the leader non-cooperative game problem can be stated only in terms of $\mathbf{a}$, and the Nash equilibrium for the leaders $\mathbf{a}^{*}\in\mathcal{A}$ can be found.
  • Follower Equilibrium Update: Using the Nash equilibrium for the leaders game problem $\mathbf{a}^{*}$, the Nash equilibrium for the follower game problem becomes $\mathbf{u}^{*}:=\mathbf{u}^{*}(\mathbf{a}^{*})$.
  • Stackelberg Equilibrium: One obtains the Class I Stackelberg equilibrium as $(\mathbf{u}^{*},\mathbf{a}^{*})\in\mathcal{U}\times\mathcal{A}$.